A can filled with water is revolved in a vertical circle of radius 4 metre and the water does not fall down. The time period of revolution will be

Grade:12th pass

1 Answers

Arun
6 years ago
 

 the top most position

 

three forces will act on the water

 

1 weight in the downward direction

 

2  normal reaction by bucket in the downward direction

 

3 centrifugal force in the upward direction

 

 

 

 so  mg + N =mv2/r

 

 for water just does not fall   N ≥ 0

 

so  mg = mv2/r

 

 v =√gr

 

 

 

 

 

so time period T = 2∏r/v

 

                         =2*3.14*4 /√4g

 

                          =3.14*4 /√9.8

 

                         =4.01 sec